OpenStudy (clairexiaoke):

-tan^2x+sec^2x i know the answer is 1 but i need to prove it

OpenStudy (clairexiaoke):

@ganeshie8 a little help please

ganeshie8 (ganeshie8):

change everything into `sin` and `cos`

OpenStudy (clairexiaoke):

\[(-\sin^2/\cos^2x)+(1/\cos^2x)\]

OpenStudy (clairexiaoke):

right this is what i got

ganeshie8 (ganeshie8):

Yes! combine the fractions

OpenStudy (clairexiaoke):

(-sin^2+1)/cos^2x

OpenStudy (clairexiaoke):

and i just get stuck here

ganeshie8 (ganeshie8):

\[\large -\dfrac{\sin^2x}{\cos^2x} + \dfrac{1}{\cos^2x}\] \[\large \dfrac{-\sin^2x+1}{\cos^2x}\]

ganeshie8 (ganeshie8):

next use below identity : \[1 = \sin^2x + \cos^2x\]

OpenStudy (clairexiaoke):

right that's the identity i thought i would need to use but i don't know how to apply it though

ganeshie8 (ganeshie8):

\[\large -\dfrac{\sin^2x}{\cos^2x} + \dfrac{1}{\cos^2x}\] \[\large \dfrac{-\sin^2x+\color{red}{1}}{\cos^2x}\] \[\large \dfrac{-\sin^2x+\color{red}{\sin^2x + \cos^2x}}{\cos^2x}\]

OpenStudy (clairexiaoke):

ok i see now so -sin^2 and sin^2x cancels out, then cos^2x/cos^2x will be one

ganeshie8 (ganeshie8):

thats it!

OpenStudy (clairexiaoke):

thank you so much! :)

ganeshie8 (ganeshie8):

np:)